# Theory of superconductivity in thin films under an external electric field

### v1

@@ Abstract
The supercurrent field effect is experimentally realized in various nano-scale devices, based on the superconductivity suppression by external electric fields being effective for confined systems. In spite of intense research, a microscopic theory and explanation of this effect is missing. Here, a microscopic theory of phonon-mediated superconductivity in thin films is presented, which accounts for the effect of quantum confinement on the electronic density of states, on the Fermi energy, and on the topology of allowed states in momentum space. By further accounting for the interplay between quantum confinement, the external static electric field, the Thomas-Fermi screening in the electron-phonon matrix element, and the effect of confinement on the Coulomb repulsion parameter, the theory predicts the critical value of the external electric field as a function of the film thickness, above which superconductivity is suppressed. In particular, this critical value of the electric field is the lower the thinner the film, in agreement with recent experimental observations. Crucially, this effect is predicted by the theory when both Thomas-Fermi screening and the Coulomb pseudopotential are taken into account, along with the respective dependence on thin film thickness. This microscopic theory of the supercurrent field-effect opens up new possibilities for electric-field gated quantum materials.
@@
(i) the Fermi sphere, as a whole, is shifted, in $k$-space, in the direction of the applied EF by a quantity $mv_{d}/\hbar$, where $v_{d}$ is the drift velocity of the conduction electrons in the applied EF and $m$ is the electron mass. This effect has no influence at all on the quantities relevant for superconductivity since the Fermi energy and the structure and topology of electron states in $k$-space are not altered;
@@
which is the inverse Thomas-Fermi screening length, with $n$ the number density of electrons. For a non-degenerate electron gas (the Debye-Hückel screening), $k_{F}=\sqrt{\frac{4\pi e^{2}n}{k_{B}T}}$, with $e$ the electron charge and $k_{B}$ the Boltzmann constant.
When screening is large compared to the momentum of the mediating phonon, $k_{s}\equiv k_{TF}\gg q$, one can approximate the above expression Eq. (9), as simply:
@@
We can thus replace this result in Eq. (7), and, in the case of large screening, obtain:
@@
represents the characteristic energy scale of phonons for pairing in the strong-coupling limit, $f_{1}$, $f_{2}$ are semi-empirical correction factors, as defined in  [30] and $u^{\star}$ is the Coulomb repulsion parameter. As a consistency check, the dependence of $\lambda$ on $N(\epsilon_{F})$ in Eq. (6) is such that, within Eq. (17), one correctly recovers the dependence of $T_{c}$ on $N(\epsilon_{F})$ as predicted by the BCS formula in the weak-coupling limit.
@@
$E_{cr}=2\Delta/e,$ (31)
where $\Delta$ is the BCS energy gap, which also represents the binding energy of a Cooper pair, and $e$ is the electron charge. Within a simplified BCS picture where screening is not taken into account, in the regime $L>L_{c}$ we can write:
@@
$E_{cr}=4\frac{\hbar\omega_{D}}{e}\left[\exp\left(\frac{1}{N^{bulk}(\epsilon_{F})\left(1+\frac{2}{3}\frac{\pi}{nL^{3}}\right)^{1/3}U}\right)-1\right]^{-1}.$ (33)
@@
$E_{cr}=4\frac{\hbar\omega_{D}}{e}\left[\exp\left(\frac{1}{2\frac{Vm\sqrt{Ln}}{\sqrt{2}\pi^{3/2}\hbar^{2}}U}\right)-1\right]^{-1}.$ (34)
@@
$E_{cr}=2\Delta/e\approx 2\cdot\frac{1.764}{e}\,k_{B}T_{c}$ (35)
@@
$E_{cr}$ $=C\exp\left[-\frac{1.04\left[1+\lambda\left(L\right)\right]}{\lambda\left(L\right)-u^{\star}(L)\left[1+0.62\lambda\left(L\right)\right]}\right],$
@@
where $C=2.94k_{B}\,f_{1}\,f_{2}\,\omega_{log}/e$. Here, $\lambda(L)$ decreases monotonically as $L$ decreases (see above), and $u^{\star}(L)$ also decreases monotonically with $L$ with basically the same law as $\lambda$ because $u^{\star}\propto N(\epsilon_{F})$. Since $u^{\star}<\lambda$, the overall trend of $E_{cr}$ with $L$ is still dominated by $\lambda(L)$. Hence, since the exponential factor in Eq. (37) is well-known to be a monotonically increasing function of $\lambda$, cfr. Ref. [30], for all known materials, the above equation implies that a lower critical value of EF to suppress superconductivity is required upon decreasing the film thickness, which is at odds with experimental observations [34].
@@
$E_{cr}$ $=C\exp\left[-\frac{1.04\left[1+\lambda\right]}{\lambda-u^{\star}(L)\left[1+0.62\lambda\right]}\right],$
@@
that is, the only dependence on the film thickness $L$ is the one contributed by the Coulomb repulsion parameter $u^{\star}(L)$ according to Eq. (30) and Eq. (19). Since $\Delta\sim E_{cr}$ is a monotonically decreasing function of $u^{\star}(L)$, and $u^{\star}(L)$ is a monotonically decreasing function of $L$, this result directly implies that $E_{cr}$, i.e. the critical EF value to suppress superconductivity, decreases upon decreasing the film thickness $L$. This is the most important finding of the present Letter, which explains the experimental observations of Ref. [34].
@@
The theory presents a basis to explain and understand the field-effect on supercurrent in thin films, in particular, to explain the suppression of superconductivity as a function of the film thickness and the relationship between the critical electric field and the film thickness. A crucial role in the ability of the theory to explain the field-effect on supercurrent in thin films is played by the synergy between the Thomas-Fermi screening in the electron-phonon matrix element, and electron-electron Coulomb repulsion. In particular, thanks to the Thomas-Fermi screening and its dependence on the thickness-dependent Fermi energy, the electron-phonon coupling $\lambda$ becomes independent of film thickness due to an exact cancellation of the opposite effects. This leaves, as the only active thickness dependence, the one of the Coulomb parameter $u^{\star}$, which increases upon decreasing the film thickness, thus lowering the energy gap as the thickness decreases. Furthermore, we have estimated the critical value of the external electric field, $E_{cr}$, required to suppress the superconductivity under these conditions. When both the Thomas-Fermi screening and the Coulomb repulsion, with their respective dependencies on the film thickness, are duly taken into account, $E_{cr}$ decreases upon decreasing the thickness of the thin film, in agreement with the known phenomenology [4, 14].
All in all, these results will pave the way for the rational and quantitative optimization of the field-effect transistors and for electric-field gated quantum materials, and it can be extended in a future work to ultrathin films of few atomic layers [37].Furthermore, a wealth of new fundamental laws relating physical quantities have been presented, which are testable in future experiments.
.1 Acknowledgments
A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”, from US Army Research Office through contract nr. W911NF-22-2-0256, and from the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. V. M. F. gratefully acknowledges support by the European Cooperation in Science and Technology via COST Action CA21144 (SUPERQUMAP) and fruitful discussions with F. Giazotto, R. Córdoba, J. M. De Teresa, A. Di Bernardo, M. Cuoco, and G. Ummarino.

### v2

@@ Abstract
The supercurrent field effect is experimentally realized in various nano-scale devices, based on the superconductivity suppression by external electric fields being effective for confined systems. In spite of intense research, a microscopic theory of this effect is missing. Here, a microscopic theory of phonon-mediated superconductivity in thin films under an external electric field is presented, which accounts for the effect of quantum confinement on the electronic density of states, on the Fermi energy, and on the electron Coulomb repulsion. By accounting for the complex interplay between quantum confinement, the external static electric field, the Thomas-Fermi screening in the electron-phonon matrix element, and the effect of confinement on the Coulomb repulsion parameter, the theory predicts the critical value of the external electric field as a function of the film thickness, above which superconductivity is suppressed. In particular, this critical value of the electric field is exponentially lower the thinner the film, in agreement with recent experimental observations. Crucially, this effect is predicted by the theory when both Thomas-Fermi screening and the Coulomb pseudopotential are taken into account, along with the respective dependence on the thin film thickness. This microscopic theory opens up new possibilities for the supercurrent field effect and for electric-field gated quantum materials.
I Introduction
@@ II.1 Eliashberg formalism
II.2 External electric field and Thomas-Fermi screening
@@
(i) the Fermi sphere, as a whole, is shifted, in $k$-space, in the direction of the applied EF by a quantity $mv_{d}/\hbar$, where $v_{d}$ is the magnitude of the drift velocity of the conduction electrons in the applied EF and $m$ is the electron mass. This effect has no influence at all on the quantities relevant for superconductivity since the Fermi energy and the structure and topology of electron states in $k$-space are not altered;
@@
which is the inverse Thomas-Fermi screening length, with $n$ the number density of electrons. For a non-degenerate electron gas (the Debye-Hückel screening), $k_{s}=\sqrt{\frac{4\pi e^{2}n}{k_{B}T}}$, with $e$ the electron charge and $k_{B}$ the Boltzmann constant.
When the screening wavevector is large compared to the momentum of the mediating phonon, $k_{s}\equiv k_{TF}\gg q$, one can approximate the above expression Eq. (9), as simply:
@@
We can thus replace this result in Eq. (7), and, in the case of strong screening, obtain:
@@ II.3 Critical superconducting temperature
represents the characteristic energy scale of phonons for pairing in the strong-coupling limit, $f_{1}$, $f_{2}$ are semi-empirical correction factors of order unity, as defined in  [30], and $u^{\star}$ is the Coulomb repulsion parameter. As a consistency check, the dependence of $\lambda$ on $N(\epsilon_{F})$ in Eq. (6) is such that, within Eq. (17), one correctly recovers the dependence of $T_{c}$ on $N(\epsilon_{F})$ as predicted by the BCS formula in the weak-coupling limit.
II.4 Confinement effects of thin film thickness
@@
III Results: thickness-dependent suppression of superconductivity
III.1 Derivation with thickness-independent Coulomb repulsion
@@ III.2 Thickness-dependent effect of Coulomb repulsion
III.3 Thickness-dependent critical electric field for superconductivity suppression
@@
$E_{cr}=\frac{2\Delta}{e\,\xi},$ (31)
where $\Delta$ is the BCS energy gap, which also represents the binding energy of a Cooper pair, $e$ is the electron charge, and $\xi$ is the coherence length, which is a measure of the spatial distance between two electrons forming a Cooper pair. Within a simplified BCS picture where screening is not taken into account, in the regime $L>L_{c}$ we can write:
@@
$E_{cr}=4\frac{\hbar\omega_{D}}{e\,\xi}\left[\exp\left(\frac{1}{N^{bulk}(\epsilon_{F})\left(1+\frac{2}{3}\frac{\pi}{nL^{3}}\right)^{1/3}U}\right)-1\right]^{-1}.$ (33)
@@
$E_{cr}=4\frac{\hbar\omega_{D}}{e\,\xi}\left[\exp\left(\frac{1}{2\frac{Vm\sqrt{Ln}}{\sqrt{2}\pi^{3/2}\hbar^{2}}U}\right)-1\right]^{-1}.$ (34)
@@
$E_{cr}=2\frac{\Delta}{e\,\xi}\approx 2\cdot\frac{1.764}{e\,\xi}\,k_{B}T_{c}$ (35)
@@
$E_{cr}$ $=\frac{C}{\xi}\exp\left[-\frac{1.04\left[1+\lambda\left(L\right)\right]}{\lambda\left(L\right)-u^{\star}(L)\left[1+0.62\lambda\left(L\right)\right]}\right],$
@@
where $C=2.94k_{B}\,f_{1}\,f_{2}\,\omega_{log}$. Here, $\lambda(L)$ decreases monotonically as $L$ decreases (see above), and $u^{\star}(L)$ also decreases monotonically with $L$ with basically the same law as $\lambda$ because $u^{\star}\propto N(\epsilon_{F})$. Since $u^{\star}<\lambda$, the overall trend of $E_{cr}$ with $L$ is still dominated by $\lambda(L)$ inside the exponential. Hence, since the exponential factor in Eq. (37) is well-known to be a monotonically increasing function of $\lambda$, cfr. Ref. [30], for all known materials, the above equation implies that a lower critical value of EF to suppress superconductivity is required upon decreasing the film thickness, which is at odds with experimental observations [4].
@@
$E_{cr}$ $=\frac{C}{\xi}\exp\left[-\frac{1.04\left[1+\lambda\right]}{\lambda-u^{\star}(L)\left[1+0.62\lambda\right]}\right],$
@@
that is, the dominant dependence on the film thickness $L$ is the one contributed by the Coulomb repulsion parameter $u^{\star}(L)$ inside the exponential factor, via Eq. (30) and Eq. (19). Since $\Delta\sim E_{cr}$ is a monotonically decreasing function of $u^{\star}(L)$, and $u^{\star}(L)$ is a monotonically decreasing function of $L$, this result directly implies that $E_{cr}$, i.e. the critical EF value to suppress superconductivity, decreases upon decreasing the film thickness $L$. This is the most important finding of the present paper, which explains some experimental observations of Ref. [4]. While in the above discussion we neglected the possible dependence of the coherence length $\xi$ on $L$, which cannot be predicted a priori since it is a function of the actual film fabrication process, we expect this dependence to be sub-dominant with respect to the $L$-dependence of the exponential factor in Eq. (38).
@@
Material and temperature dependence of the critical field To summarize, the main result of our theory is Eq. (38) with the thickness-dependent Coulomb parameter given by Eq. (30) evaluated with Eq. (19).
The dependence of the Coulomb parameter $u^{\star}$ on the film thickness $L$ controls the thickness-dependence of the critical field $E_{cr}$. This is a universal dependence on $L$, valid for all materials and independent of composition.
In the above equation for the critical EF, a possibly material-dependent parameter is the coherence length $\xi$. The latter is given by [35]:
$\frac{1}{\xi}=\frac{1}{\xi_{0}}+\frac{1}{\ell}$ (39)
where $\xi_{0}$ is the intrinsic (Pippard) coherence length, and $\ell$ is the mean free path. Thin films, such as those used in the supercurrent field effect devices, have a microstructure characterized by microcrystallites, the size of which sets the value of $\ell$. Since, typically, $\ell\ll\xi_{0}$ (because $\xi_{0}$ can be tens or hundreds of nanometers), the coherence length $\xi$ is controlled by $\ell$, and, hence, by the fabrication process.
The experimental observation that $E_{cr}$ is within two orders of magnitude for different materials, as shown in Fig. 3 of Ref. [36], is explained by the fact that $E_{cr}\sim\Delta(L)/\xi$, where $\Delta(L)$ is the thickness-dependent energy gap. Because the thickness dependence of $\Delta(L)$ is the same for all materials, the only material dependence is that of the energy gap of the bulk material. Since (i) $\Delta\simeq 1.76k_{B}T_{c}$ for the bulk material, where $T_{c}$ varies for conventional superconductors within a factor of 20 [37], and (ii) the coherence length $\xi$ in thin films, which is set by $\ell$, is estimated to vary by at most an order of magnitude (1-10 nm), depending on the fabrication process, it is clear that $E_{cr}$ can vary within two orders of magnitude.
We also notice that, as a matter of fact, the above thickness dependence of the critical electric field $E_{cr}$ is the same at all temperatures. This is evident from the above Eq. (38), and from the fact that our confinement model for the quasiparticles is entirely temperature-independent (since it is a purely quantum effect).
Finally, since the coherence length $\xi$ is a measure of the physical separation between two electrons forming a Cooper pair, it is clear that the above theory is applicable only to systems where $\xi<L$.
IV Conclusions
@@
The theory presents a possible basis to qualitatively explain and understand the field-effect on supercurrent in thin films. In particular, to explain the suppression of superconductivity as a function of the film thickness and the relationship between the critical electric field and the film thickness. A crucial role in the ability of the theory to explain the thickness-dependent suppression of superconductivity in thin films is played by the synergy between the Thomas-Fermi screening in the electron-phonon matrix element, and electron-electron Coulomb repulsion. In particular, thanks to the Thomas-Fermi screening and its dependence on the thickness-dependent Fermi energy, the electron-phonon coupling $\lambda$ becomes independent of film thickness due to an exact cancellation of the opposite effects. This leaves, as the only active thickness dependence, the one of the Coulomb parameter $u^{\star}$, which increases upon decreasing the film thickness, thus lowering the energy gap as the thickness decreases. Furthermore, we have estimated the critical value of the external electric field, $E_{cr}$, required to suppress the superconductivity under these conditions. When both the Thomas-Fermi screening and the Coulomb repulsion, with their respective dependencies on the film thickness, are duly taken into account, $E_{cr}$ decreases upon decreasing the thickness of the thin film, in agreement with the known phenomenology [4, 14].
All in all, these results represent a step forward for the rational and quantitative optimization of the field-effect transistors and for electric-field gated quantum materials, and it can be extended in a future work to ultrathin films of few atomic layers [39]. Furthermore, a wealth of new fundamental laws relating physical quantities have been presented, which are testable in future experiments.
Acknowledgments
A.Z. gratefully acknowledges funding from the European Union through Horizon Europe ERC Grant number: 101043968 “Multimech”, from US Army Research Office through contract nr. W911NF-22-2-0256, and from the Niedersächsische Akademie der Wissenschaften zu Göttingen in the frame of the Gauss Professorship program. V. M. F. gratefully acknowledges support by the European Cooperation in Science and Technology via COST Action CA21144 (SUPERQUMAP) and fruitful discussions with F. Giazotto, R. Córdoba, J. M. De Teresa, A. Di Bernardo, M. Cuoco, and G. Ummarino, L. Martinoia, A. Amoretti.